Home‎ > ‎Reports‎ > ‎Crew099‎ > ‎

day06


Crew 99 Log Book for February 3, 2011

Photos of the day





Commander's Check-In Report

Donna Viola Reporting

Crew Physical Status: Good


Time Departed/Returned from EVA:

EVA #8 11:00-12:00; EVA #7 14:30-16:30


Brief Narrative of Field Mission Results: EVA planner test; endolith sample collection


EVA Data/Interpretations: as per EVA reports


Engineering/Hab Maintenance: as per engineering report; outside water tank empty, pipes freezing (thawed by afternoon), plumbing problems - toilet flushes slowly, needs frequent plunging


Report Transmission Schedule: Engineering, EVA #8, EVA #9, Chef's, Commander's. By 8:30pm


Plans for Tomorrow: to be decided


Inventory: as requested, will do inventory of remaining dehydrated

vegetables within the next few days


Support Requested: outside water tank is empty; suggestions for

plumbing problems welcome


Miscellaneous:



Commander's Report

Donna Viola Reporting

It was another long night, as Crew 99 took shifts checking on the LAMBDA experiment every hour - a ritual we will need to continue day and night until the completion of the experiment. Hence, everyone woke up a little sleepy.

However, the mission must go on - Heidi and Andie set off for the Mellow Mushrooms in the morning for another test of the EVA planner. Max spent the day in the Hab, preparing for his balloon launch project, which is slated for launch some time next week. Sukrit also opted to remain in the Hab to work with the astronomy software and CCD camera.

In the afternoon, Andie, Kevin, and I rode the ATVs south to collect some samples of endolithic colonies - and also saw plenty of lichen and cyanobacteria living happily on the rocks! We then climbed to the top of "the Pimple" to get a great view of the surrounding area.

Dinner tonight was"Forever Young Macaroni and Cheese" - so hopefully, we can finally get the song "Forever Young" out of our heads (where it's been stuck ever since we realized the name of the one of our rehydrate-only meals).

Tonight, the crew plans on relaxing and watching a sci-fi classic: Invaders from Mars, anyone? (:


Engineering  Report

Kevin Newman Reporting


Generator/Electricity:

· Kitty (Diesel Generator) (Used? Yes/No): Yes

· Honey (Gas Generator) (Used? Yes/No): No

· Battery Charge Level:

o 10.59 V everstart

o 12.37 V armor plate

o 13.32 V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):
1 generator run, all is well


Fuel Status:

· Diesel (Full/0.75/0.5/0.25/Empty):2/7

· Propane Tank (Full/0.75/0.5/0.25/Empty):77%

· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 10 Gallons

Notes/Comments:


Oil:

· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

· Honey Oil Quantity (# of Quarts in storage): Leave it to DG

· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

· Car Oil Quantity (# of Quarts in storage):


Water Status:

· Outside Potable Water Tank Level (inches from bottom): 16

· Trailer Potable Water Tank Level (inches from bottom): EMPTY

· Hab Potable Water Tank Level (inches from bottom): 11

· Potable Water Meter Reading: 416303


Notes/Comments:
Trailer tank has 1’’ layer of ice suspended 8’’ from bottom and ~1’’ of water/ice on the bottom, not enough to continue using pump to Hab potable water tank.


GreenHab:

· Condition of plants in Tank 1: Nothing

· Condition of plants in Tank 2: Ok

· Condition of plants in Tank 3: Good

· Greenhab notes/ comments:


Observatory:

· Telescope Used? (Yes/no) Yes

Observatory Notes:

Transportation:


Hab Car (NOT RENTAL)

· Used (yes/no):No

· Oil Checked (yes/no): No

· Tire Status: Not checked

· Odometer at end of day: 123597.9

· Notes/Comments on Hab Car:


ATV 1 (Opportunity):

· Used: Yes

· Oil Checked: No

· Fuel Consumed: Replaced

· Tire Status: Checked, Filled

· Comments:


ATV 2 (Spirit):

· Used: Yes

· Oil Checked: No

· Fuel Consumed: Replaced

· Tire Status: Checked, Filled

· Comments:


ATV 3 (Viking I):

· Used: No

· Oil Checked: No

· Fuel Consumed: None

· Tire Status: Not Checked

· Comments:


ATV 4 (Viking II):

· Used: Yes

· Oil Checked: No

· Fuel Consumed: Replaced

· Tire Status: Checked, Filled

· Comments:


Heating and Ventilation:

· Thermostat setting upstairs (Farenheit): 70

· Actual temperature upstairs (Farenheit): 77

· Thermostat setting downstairs (Farenheit): 65

· Actual temperature downstairs (Farenheit): 64


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: All good


General Engineering

1. Notes/Comments: Its still really cold!!!


EVA 8 Report

Andie Gilkey Reporting

Time: 11:00-12:00


Crew Members: Andie Gilkey, Heidi Beemer


Site Location:  UTM: 518800 4250250

 Transit Mode: foot


Objectives: Measure how accurately the mission planner predicts astronaut energy consumption, distance traveled, and EVA time on the way to Mellow Mushrooms.


Results: We were able to successfully measure astronaut heart rate throughout the length of the mission, as well as distance traveled and EVA time.


Lessons Learned: Use a GPS on BOTH legs of the EVA, no matter how short they are. Don't rely on your sense of direction to get back, even if it's only around the corner, otherwise you might end up taking the long way home. :-)



EVA 9 Report

Kevin Newman Reporting

Time: 14:30-16:30

Crew Members: Andie Gilkey, Donna Viola, Kevin Newman

Site Location:  N 38.39128, W 100.78496 and N 38.39417, W 110.78454 and N 38.39620 W 110.78907

 Transit Mode: ATV / Foot

Objectives: Find endoliths

Results: Found endoliths, brought back samples for future analysis, took pictures.

Lessons Learned: It is very difficult to put the ATV helmet straps back on while wearing gloves.



Journalist Report

Sukrit Ranjan Reporting


Today saw the continued testing of our MFC apparatus. After waking up in shifts all night long to agitate the MFC, the crew woke up tired. Early in the morning, Heidi and Andie left for an EVA to test the mission planner software.

In the afternoon, Andie, Kevin and Donna went on EVA to look for endoliths. Meanwhile, in the hab Sukrit and Max made progress on the observatory, imaging with the CCD for the first time! They look forward to, eventually, taking beautiful photos from Musk Observatory.

For dinner, the crew sat down to a long-awaited Forever Young mac and cheese dinner, which we all overindulged on. Tonight, we hope to sit down and relax with a movie: perhaps, "Invaders from Mars"?


Cook's Report

Kevin Newman Reporting

1a) Today was

-A non-cooking day.

1b) Was today a special day celebrated at the main meal with special food or activity?

- Of course! Today is national carrot day, and we rejoiced accordingly.

http://blog.yestocarrots.com/2009/02/its-national-carrot-day.html

2)  List the foods served at today's main meal (usually the evening meal), giving full names of each.

- Forever Young Macaroni and Cheese (it contained carrots)


3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)  (Example:  2 servings potato soup stored; 2 servings rice burned/inedible, discarded)

- None, we ate it all.

4)  Recipes for anything more complicated than following package directions exactly, or rehydrating.

- None

5)  Comments/questions for Kim and Jean, the food study investigators

- The only meal we've given a '9' for overall acceptability so far! The  macaroni and cheese was a huge hit!